OpenStudy (zzr0ck3r):

I would like to show that if \(A\) is a positive semidefinite matrix and \(P\) is a permutation matrix then \(P^TAP\) is also PSD.

OpenStudy (bobo-i-bo):

According to wikipedia, a matrix M being positive-definite is equivalent to M having all positive eigenvalues. So assuming that PSD is equivalent to non-negative eigenvalues, then it's very obvious that \(P^TAP\) is also PSD.

OpenStudy (bobo-i-bo):

My assumption is correct :) PSD matrix is equivalent to a having only non-negative eigenvalues.

OpenStudy (michele_laino):

here is my reasoning: I consider a permutation Matrix \(P\), so, from the general theory on matrices and determinants, we have: \[\det P = {\left( { - 1} \right)^n}\left( {\det {I_n}} \right) = {\left( { - 1} \right)^n},\quad {P^T}P = {I_n}\] wherein \(I_n\) is the identity Matrix, and \(n\) is the dimension, over the set of real numbers, of the involved vector space. Next we have: \[\begin{gathered} \det \left( {{P^T}AP - \lambda I} \right) = \det \left( {{P^T}AP - {P^T}\lambda IP} \right) = \hfill \\ \hfill \\ = \det \left( {{P^T}\left( {A - \lambda I} \right)P} \right) = \left( {\det {P^T}} \right)\left\{ {\det \left( {A - \lambda I} \right)} \right\}\det P = \hfill \\ \hfill \\ = {\left( {{{\left( { - 1} \right)}^n}} \right)^2}\left\{ {\det \left( {A - \lambda I} \right)} \right\} = \det \left( {A - \lambda I} \right) \hfill \\ \end{gathered} \] namely the matrices \(P^TAP\) and \(A\) have the same characteristic polynomial, and then have the same eigenvalues, so if \(A\) is positive semidefinite, als \(P^TAP\) is positive semidefinite

OpenStudy (michele_laino):

wherein the eponent \(n\) in \((-1)^n\) is the number of swaps of columns starting from the identity matrix